# Estimating non-linear functionals of trawl processes

### v2

@@ Appendix A Computing the AVAR
$F_{n,2}^{(1,1)}(s,r)=a_{n}(0)\int_{s\lor r-r\land s}^{s\lor r}a(u)\mathrm{d}u\left(1-\frac{s\lor r}{n\Delta_{n}}\right).$
Thus, in view that $\frac{\tau_{n}(r)\lor\tau_{n}(s)}{n\Delta_{n}}\leq 1$ for $r,s\leq n\Delta_{n}$ and $a_{n}\leq a\leq a(0)$, we easily conclude that $\rvert F_{n,1}^{(1,1)}(s,r)\rvert\leq 2\mathfrak{K}_{4}a(0)$ as well as $\rvert F_{n,2}^{(1,1)}(s,r)\rvert\leq 2a(0)\mathrm{Leb}(A).$ This shows that the function $F_{n}^{(1,1)}:=F_{n,1}^{(1,1)}+F_{n,2}^{(1,1)}$ is bounded and for every given $s,r\geq 0$
$F_{n}^{(1,1)}(\tau_{n}(s),\tau_{n}(r))$ $\rightarrow\mathfrak{K}_{4}a(s\lor r)+a(0)\int_{s\lor r-s\land r}^{s\lor r}a(u)\mathrm{d}u=:F^{(1,1)}(s,r),$
just as required.
@@
$\Sigma_{n}^{(2,2)}(l,l^{\prime})=$ $a_{n}(0)\sum_{j=l\lor l^{\prime}+1}^{n}\sum_{k=l\lor l^{\prime}}^{j-1}\mathrm{Leb}(\cup_{i=0}^{k-l\lor l^{\prime}}\mathcal{\mathcal{P}}_{A}^{n}(i,j))$
$=$ $a_{n}(0)\frac{1}{n}\sum_{j=l\lor l^{\prime}+1}^{n}\int_{t_{l+1}\lor t_{l^{\prime}+1}}^{t_{j+1}}a(u)\mathrm{d}u.$
Thus, (74) holds in this situation with
@@
Clearly $\rvert F_{n}^{(2,2)}(s,r)\rvert\leq a(0)\mathrm{Leb}(A)$ and by a simple change of variables and the DCT we further have that $F_{n}^{(2,2)}(\tau_{n}(s),\tau_{n}(r))\rightarrow a(0)\int_{s\lor r}^{\infty}a(u)\mathrm{d}u=:F^{(2,2)}(s,r),$ as needed.
@@
$\Sigma_{n}^{(3,3)}(l,l^{\prime})=\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}+1}^{n}\sum_{k=l\lor l^{\prime}}^{j-1}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k+1,j))\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{j}-t_{k}+t_{l}\lor t_{l^{\prime}}}a(u)\mathrm{d}u,$
By letting $m=j-k+l\lor l^{\prime}$ and denoting $A_{m}^{l,l^{\prime}}:=\int_{t_{m-1}-t_{l}\lor t_{l^{\prime}}}^{t_{m}-t_{l}\lor t_{l^{\prime}}}a(s)\mathrm{d}s$, we deduce that
$\sum_{k=l\lor l^{\prime}}^{j-1}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k+1,j))\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{j}-t_{k}+t_{l}\lor t_{l^{\prime}}}a(s)\mathrm{d}s$ $=\sum_{m=l\lor l^{\prime}+1}^{j}\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{m}}a(u)\mathrm{d}u[A_{m}^{l,l^{\prime}}-A_{m+1}^{l,l^{\prime}}]$
$=\Delta_{n}a_{n}(0)\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{l+1}\lor t_{l^{\prime}+1}}a(u)\mathrm{d}u$
$+\sum_{m=l\lor l^{\prime}+2}^{j}\int_{t_{m-1}}^{t_{m}}a(u)\mathrm{d}u\int_{t_{m-1}}^{t_{m}}a(u-t_{l}\lor t_{l^{\prime}})\mathrm{d}u$
$-\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{j}}a(u)\mathrm{d}u\int_{t_{j}}^{t_{j+1}}a(u-t_{l}\lor t_{l^{\prime}})\mathrm{d}u$
$=\Delta_{n}a_{n}(0)\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{l+1}\lor t_{l^{\prime}+1}}a(u)\mathrm{d}u$
@@
$+\mathrm{O}\left(\int_{t_{j}}^{t_{j+1}}a(u-t_{l}\lor t_{l^{\prime}})\mathrm{d}u\right)$
In consequence, we have that $\Sigma_{n}^{(3,3)}(l,l^{\prime})=F_{n}^{(3,3)}(t_{l},t_{l^{\prime}})+\mathrm{O}(1/n\Delta_{n}),$ where
$F_{n}^{(3,3)}(s,r):=\frac{1}{n\Delta_{n}}\int_{s\lor r}^{n\Delta_{n}}\int_{s\lor r}^{y}a_{n}(\tau_{n}(u))a(u-s\lor r)\mathrm{d}u\mathrm{d}y+a(0)\int_{s\lor r-s\land r}^{s\lor r}a(u)\mathrm{d}u\left(1-\frac{s\lor r}{n\Delta_{n}}\right).$
@@
$F_{n}^{(3,3)}(\tau_{n}(s),\tau_{n}(r))\rightarrow\int_{s\lor r}^{\infty}a(u)a(u-s\lor r)\mathrm{d}u+a(0)\int_{s\lor r-r\land s}^{s\lor r}a(u)\mathrm{d}u=:F^{(3,3)}(s,r).$
@@
Repeating the same reasoning as above we get in this case that
$\Sigma_{n}^{(4,4)}(l,l^{\prime})$ $=\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}+1}^{n}\sum_{k=l\lor l^{\prime}}^{j-1}\mathrm{Leb}(\cup_{m=k+1}^{j}\mathcal{\mathcal{P}}_{A}^{n}(k+1,m))\mathrm{Leb}(\cup_{i=0}^{k-l\lor l^{\prime}}\mathcal{P}_{A}^{n}(i,j))$
$=\frac{1}{n}\sum_{j=l\lor l^{\prime}+1}^{n}\sum_{k=l\lor l^{\prime}}^{j-1}\int_{t_{j}-t_{k}}^{t_{j+1}-t_{k}}[a_{n}(0)-a(u)]a_{n}(t_{j}-t_{k}+t_{l}\lor t_{l^{\prime}})\mathrm{d}u$
$=\frac{1}{n}\sum_{j=l\lor l^{\prime}+1}^{n}\sum_{m=l\lor l^{\prime}+1}^{j-1}\int_{t_{m}-t_{l}\lor t_{l^{\prime}}}^{t_{m+1}-t_{l}\lor t_{l^{\prime}}}[a_{n}(0)-a(u)]a_{n}(t_{m})\mathrm{d}u$
$=F_{n}^{(4,4)}(t_{l},t_{l^{\prime}})+\mathrm{O}(\Delta_{n})$
with
@@
It is not difficult to see that $F_{n}^{(4,4)}$ is bounded by $2a(0)(\Delta_{n}a(0)+\mathrm{Leb}(A))$ and arguing as before we also have that
$F_{n}^{(4,4)}(\tau_{n}(s),\tau_{n}(r))\rightarrow\int_{s\lor r}^{\infty}[a(0)-a(u-s\lor r)]a(u)\mathrm{d}u=:F^{(4,4)}(s,r).$
@@
Hence, $\Sigma_{n}^{(1,2)}(l,l^{\prime})=F_{n}^{(1,2)}(t_{l},t_{l^{\prime}})+\mathrm{O}(1/n),$ in which
$F_{n}^{(1,2)}(s,r):=\frac{1}{n}\sum_{j=[s/\Delta_{n}]\lor[r/\Delta_{n}]}^{n}\sum_{m=[r/\Delta_{n}]+1}^{j-(j-[s/\Delta_{n}]-[r/\Delta_{n}])^{+}}\int_{t_{m}}^{t_{m+1}}a(u)a_{n}(s+r-\tau_{n}(u))\mathrm{d}u$
which is obviously bounded by $a(0)\mathrm{Leb}(A)$ and for $n$ large enough and fix $s,r\geq 0$ we further have
@@
Using that
$\mathbb{E}(\beta_{j,j-1,l}^{(2)}\beta_{k,j-1,l^{\prime}}^{(2)})=\begin{cases}\sum_{i=0}^{k-l^{\prime}}\mathrm{Leb}(\cup_{m=j-l}^{j-1}\mathcal{\mathcal{P}}_{A}^{n}(i,m)),&\text{if }k\leq j-(l-l^{\prime});\\ \sum_{m=k-l^{\prime}}^{j-1}\mathrm{Leb}(\cup_{i=0}^{j-l}\mathcal{\mathcal{P}}_{A}^{n}(i,m)),&\text{otherwise},\end{cases}$
we decompose
@@
where
@@
$\mathfrak{S}_{j,l,l^{\prime}}^{n,1}=$ $\Delta_{n}\sum_{m=l\lor(l^{\prime}+1)}^{j-1}\int_{t_{m}}^{t_{m+1}}[a(u-t_{l})-a(u)][a_{n}((t_{l}-t_{l^{\prime}+1})^{+})-a_{n}(t_{m}-t_{l^{\prime}})]\mathrm{d}u$
@@
Hence, (74) holds with
$-F_{n}^{(1,3)}(s,r)$ $=\frac{1}{n\Delta_{n}}\int_{s\lor r}^{n\Delta_{n}}\int_{s\lor r}^{y}[a(u-s)-a(u)][a_{n}((s-r-\Delta_{n})^{+})-a_{n}(\tau_{n}(u)-s)]\mathrm{d}u\mathrm{d}y$
$+\int_{0}^{s}a(u)[a_{n}((s-r-\Delta_{n}-\tau_{n}(u))^{+})-a_{n}((s-r-\Delta_{n})^{+})]\mathrm{d}u\left(1-\frac{s\lor r}{n\Delta_{n}}\right),\,\,$
for $s,r\leq n\Delta_{n},$ and by the arguments used before satisfies that $\rvert F_{n}^{(1,3)}(s,r)\rvert\leq 4a(0)\mathrm{Leb}(A)$ and
@@
$\mathbb{E}(\beta_{j,j,l}^{(1)}\chi_{j}\alpha_{k+1,j}\beta_{k,j,l^{\prime}}^{(1)})$ $=\begin{cases}2\mathbb{E}(\beta_{k,l^{\prime},j}^{(1)2})\mathbb{E}(\alpha_{k+1,j}^{2})&\text{if }k\leq j-l-1;\\ \mathbb{E}(\beta_{k,l^{\prime},j}^{(1)}\beta_{j,l,j}^{(1)})\mathbb{E}(\alpha_{k+1,j}^{2})&\text{if }k\geq j-l;\end{cases}$
$=\mathbb{E}(\beta_{k,l^{\prime},j}^{(1)}\beta_{j,l,j}^{(1)})\mathbb{E}(\alpha_{k+1,j}^{2})(1+\mathbf{1}_{k\leq j-l-1}).$
@@
$\sum_{k=l^{\prime}}^{j-1}\mathbb{E}(\varsigma_{k,j}\beta_{j,j-1,l}^{(2)})\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,j))=$ $\Delta_{n}\sum_{k=l^{\prime}}^{j-l-1}\int_{t_{j-k-1}}^{t_{j-k}}\left[a(u-t_{l})-a(u)\right]\mathrm{d}ua_{n}(t_{j-k-1}+t_{l^{\prime}+1})$
$=$ $\Delta_{n}\sum_{m=l}^{j-l^{\prime}-1}\int_{t_{m}}^{t_{m+1}}\left[a(u-t_{l})-a(u)\right]\mathrm{d}ua_{n}(t_{m}+t_{l^{\prime}+1}).$
Hence, $\Sigma_{n}^{(1,4)}(l,l^{\prime})=F_{n}^{(1,4)}(t_{l},t_{l^{\prime}})+\mathrm{O}(\Delta_{n}),$ with
@@
Obviously, $\rvert F_{n}^{(1,4)}(s,r)\rvert\leq 2a(0)\mathrm{Leb}(A)$ and by the DCT we also get that
@@
$-\Sigma_{n}^{(2,3)}(l,l^{\prime})=$ $\sum_{j=l\lor l^{\prime}+1}^{n}\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{j-1}\mathbb{E}(\chi_{k}\beta_{k^{\prime},j-1,l^{\prime}}^{(2)})\mathbb{E}(\alpha_{k^{\prime}+1,j}^{2})\mathbf{1}_{k^{\prime}\leq(k-l-1)\land(k+l^{\prime})}\mathbf{1}_{k\geq l+l^{\prime}+1}\mathbf{1}_{j\geq l^{\prime}+l^{\prime}+2}$
$=$ $\frac{1}{n\Delta_{n}}\sum_{k=l+l^{\prime}+1}^{n-1}\sum_{k^{\prime}=l^{\prime}}^{k-l-1}\mathrm{Leb}(\cup_{i=0}^{k^{\prime}-l^{\prime}}\mathcal{\mathcal{P}}_{A}^{n}(i,k))\sum_{j=k+1}^{n}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k^{\prime}+1,j))$
@@
which is bounded by $a(0)\mathrm{Leb}(A)$ and satisfies that
@@
Similarly as above we get that
$-\Sigma_{n}^{(2,4)}(l,l^{\prime})=$ $\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}+1}^{n-1}\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{j-1}\mathbb{E}(\chi_{k}\varsigma_{k^{\prime},j})\mathbb{E}(\beta_{k,j,l}^{(1)}\beta_{k^{\prime},j,l^{\prime}}^{(1)}).$
In this situation, after some simple change of variables, we infer that
$\sum_{k=l,k^{\prime}=l^{\prime}}^{j-1}\mathbb{E}(\chi_{k}\varsigma_{k^{\prime},j})\mathbb{E}(\beta_{k,j,l}^{(1)}\beta_{k^{\prime},j,l^{\prime}}^{(1)})=$ $\sum_{k=l,k^{\prime}=l^{\prime}}^{j-1}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k^{\prime}+1,k))\mathrm{Leb}(\cup_{i=0}^{(k^{\prime}-l^{\prime})\wedge(k-l)}\mathcal{\mathcal{P}}_{A}^{n}(i,j))\mathbf{1}_{l^{\prime}\leq k^{\prime}\leq k-1}$
$=$ $\sum_{k=l,k^{\prime}=l^{\prime}}^{j-1}\mathrm{Leb}(\cup_{i=0}^{k-l}\mathcal{\mathcal{P}}_{A}^{n}(i,j))\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k^{\prime}+1,k))\mathbf{1}_{k-1\geq k^{\prime}\geq k-(l-l^{\prime})}$
@@
Therefore, in situation we may choose
$-F^{(2,4)}(s,r)=$ $[a_{n}(0)-a_{n}((s-r)^{+})]\frac{1}{n\Delta_{n}}\int_{s\lor r}^{n\Delta_{n}}\int_{s}^{y}a(u)\mathrm{d}u\mathrm{d}y$
$+\frac{1}{n\Delta_{n}}\int_{s\lor r}^{n\Delta_{n}}\int_{s\lor r}^{y}a(u)[a_{n}((s-r)^{+})-a_{n}(\tau_{n}(u)-r-\Delta_{n})]\mathrm{d}u\mathrm{d}y,$
@@
$F^{(2,4)}(s,r)=[a(0)-a((s-r)^{+})]\int_{s}^{\infty}a(u)\mathrm{d}u\mathrm{d}y+\int_{s\lor r}^{\infty}a(u)[a((s-r)^{+})-a(u-r)]\mathrm{d}u\mathrm{d}y.$
@@
Finally, in view that for all $l,l^{\prime}\geq 0$, $k\geq l$ and $k^{\prime}\geq l^{\prime}$
$\mathbb{E}[\beta_{k,j-1,l}^{(2)}\varsigma_{k^{\prime},j}]\mathbb{E}[\alpha_{k+1,j}\beta_{k^{\prime},j,l^{\prime}}^{(1)}]\mathbf{1}_{k+1\leq k^{\prime}-l^{\prime}}=\mathbb{E}[\beta_{k,j-1,l}^{(2)}\varsigma_{k^{\prime},j}]\mathbb{E}[\alpha_{k+1,j}^{2}]\mathbf{1}_{k^{\prime}+l+1\leq k\leq k^{\prime}-l^{\prime}-1}=0,$
we conclude that in this situation $\Sigma_{n}^{(3,4)}(l,l^{\prime})=0$.
We are left to show that $\sum_{\ell,\ell^{\prime}=1}^{4}F_{n}^{(\ell,\ell^{\prime})}(\tau_{n}(s),\tau_{n}(r))\rightarrow\sum_{\ell,\ell^{\prime}=1}^{4}F^{(\ell,\ell^{\prime})}(s,r)=\Sigma_{a}(s,r).$ To see this, first note that due to the fact that $\Sigma_{n}^{(\ell^{\prime},\ell)}(l,l^{\prime})=\Sigma_{n}^{(\ell,\ell^{\prime})}(l^{\prime},l)$ and the preceding computations, we have that
$\sum_{\ell,\ell^{\prime}=1}^{4}F^{(\ell,\ell^{\prime})}(s,r)=\sum_{\ell=1}^{4}F^{(\ell,\ell)}(s,r)+\sum_{\ell=1}^{3}\sum_{\ell^{\prime}=\ell+1}^{4}\widetilde{F^{(\ell,\ell^{\prime})}(s,r)}$.
Simple but tedious computations now gives that
$\sum_{\ell=1}^{4}F^{(\ell,\ell^{\prime})}(s,r)=\mathfrak{K}_{4}a(s\lor r)+2a(0)\int_{\rvert s-r\rvert}^{\infty}a(u)\mathrm{d}u,$
@@
$\sum_{\ell=1}^{3}\sum_{\ell^{\prime}=\ell+1}^{4}F^{(\ell,\ell^{\prime})}(s,r)$ $=\int_{(s-r)^{+}}^{\infty}a(u+r-s)a(u)\mathrm{d}u+a((s-r)^{+})\int_{0}^{(r-s)^{+}}a(u)\mathrm{d}u$
$-\int_{0}^{\infty}a(u)a((s-r-u)^{+}))\mathrm{d}u+\int_{0}^{\infty}a(u+r)a(\rvert s-u\rvert)\mathrm{sgn}(s-u)\mathrm{d}u$
The sought-after identity follows from this and simple algebraic manipulations.∎

### v3

@@ Appendix A Approximations of the Asymptotic Variance
$F_{n,2}^{(1,1)}(s,r)=a_{n}(0)\int_{|s-r|}^{s\lor r}a(u)\,\mathrm{d}u\Big(1-\frac{s\lor r}{n\Delta_{n}}\Big).$
Since $a_{n}\leq a\leq a(0)$, $F_{n}^{(1,1)}:=F_{n,1}^{(1,1)}+F_{n,2}^{(1,1)}$ is bounded. Moreover,
$F_{n}^{(1,1)}(\tau_{n}(s),\tau_{n}(r))\to\mathfrak{K}_{4}a(s\lor r)+a(0)\int_{|s-r|}^{s\lor r}a(u)\,\mathrm{d}u=:F^{(1,1)}(s,r).$
@@
$\Sigma_{n}^{(2,2)}(l,l^{\prime})=a_{n}(0)\frac{1}{n}\sum_{j=l\lor l^{\prime}+1}^{n}\int_{t_{l+1}\lor t_{l^{\prime}+1}}^{t_{j+1}}a(u)\mathrm{d}u.$
Thus, (75) holds with
@@
Furthermore, by the DCT
$F_{n}^{(2,2)}(\tau_{n}(s),\tau_{n}(r))\rightarrow a(0)\int_{s\lor r}^{\infty}a(u)\mathrm{d}u=:F^{(2,2)}(s,r)$
@@
$\Sigma_{n}^{(3,3)}(l,l^{\prime})=\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}+1}^{n}\sum_{k=l\lor l^{\prime}}^{j-1}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k+1,j))\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{j}-t_{k}+t_{l}\lor t_{l^{\prime}}}a(u)\mathrm{d}u,$ (78)
For $m\geq l\lor l^{\prime}+1$, set $A_{m}^{l,l^{\prime}}:=\int_{t_{m-1}-t_{l}\lor t_{l^{\prime}}}^{t_{m}-t_{l}\lor t_{l^{\prime}}}a(s)\mathrm{d}s$, and let $m=j-k+l\lor l^{\prime}$. Then the inner sum in (78) equals
$\sum_{m=l\lor l^{\prime}+1}^{j}\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{m}}a(u)\mathrm{d}u[A_{m}^{l,l^{\prime}}-A_{m+1}^{l,l^{\prime}}]=$ $\Delta_{n}a_{n}(0)\int_{t_{l}\lor t_{l^{\prime}}-t_{l}\land t_{l^{\prime}}}^{t_{l+1}\lor t_{l^{\prime}+1}}a(u)\mathrm{d}u$
@@
$+\mathrm{O}\left(\int_{t_{j}}^{t_{j+1}}a(u-t_{l}\lor t_{l^{\prime}})\mathrm{d}u\right).$
Hence, $\Sigma_{n}^{(3,3)}(l,l^{\prime})=F_{n}^{(3,3)}(t_{l},t_{l^{\prime}})+\mathrm{O}(1/(n\Delta_{n})),$ where
$F_{n}^{(3,3)}(s,r):=$ $\frac{1}{n\Delta_{n}}\int_{s\lor r}^{n\Delta_{n}}\int_{s\lor r}^{y}a_{n}(\tau_{n}(u))a(u-s\lor r)\mathrm{d}u\mathrm{d}y$
$+a(0)\int_{s\lor r-s\land r}^{s\lor r}a(u)\mathrm{d}u\left(1-\frac{s\lor r}{n\Delta_{n}}\right).$
@@
$F_{n}^{(3,3)}(\tau_{n}(s),\tau_{n}(r))\rightarrow\int_{s\lor r}^{\infty}a(u)a(u-s\lor r)\mathrm{d}u+a(0)\int_{\rvert s-r\rvert}^{s\lor r}a(u)\mathrm{d}u=:F^{(3,3)}(s,r).$
@@
Using (63) and making a change of variable, we obtain
$\Sigma_{n}^{(4,4)}(l,l^{\prime})=\frac{1}{n}\sum_{j=l\lor l^{\prime}+1}^{n}\sum_{m=l\lor l^{\prime}+1}^{j-1}\int_{t_{m}-t_{l}\lor t_{l^{\prime}}}^{t_{m+1}-t_{l}\lor t_{l^{\prime}}}[a_{n}(0)-a(u)]a_{n}(t_{m})\mathrm{d}u.$
As a result, (75) is satisfied with
@@
In addition,
$F_{n}^{(4,4)}(\tau_{n}(s),\tau_{n}(r))\rightarrow\int_{s\lor r}^{\infty}[a(0)-a(u-s\lor r)]a(u)\mathrm{d}u=:F^{(4,4)}(s,r),$
thanks to the DCT.
@@
Hence, $\Sigma_{n}^{(1,2)}(l,l^{\prime})=F_{n}^{(1,2)}(t_{l},t_{l^{\prime}})+\mathrm{O}(1/n),$ where
$F_{n}^{(1,2)}(s,r):=\frac{1}{n}\sum_{j=[s/\Delta_{n}]\lor[r/\Delta_{n}]}^{n}\sum_{m=[r/\Delta_{n}]+1}^{j-(j-[s/\Delta_{n}]-[r/\Delta_{n}])^{+}}\int_{t_{m}}^{t_{m+1}}a(u)a_{n}(s+r-\tau_{n}(u))\mathrm{d}u.$
The bound $a_{n}\leq a\leq a(0)$ yields boundedness of $F_{n}^{(1,2)}$. For fixed $s,r\geq 0$, we further have
@@
Write
@@
where $\mathfrak{S}_{j,l,l^{\prime}}^{n,1}$ and $\mathfrak{S}_{j,l,l^{\prime}}^{n,2}$ correspond to the cases $k\leq j-(l-l^{\prime})$ and $k>j-(l-l^{\prime})$, respectively. where
@@
$\mathfrak{S}_{j,l,l^{\prime}}^{n,1}=$ $\sum_{i=0}^{j-(l^{\prime}+1)\lor l}\mathrm{Leb}(\cup_{m=j-l}^{j-1}\mathcal{\mathcal{P}}_{A}^{n}(i,m))\sum_{k=i+l^{\prime}}^{j-1-(l-l^{\prime}+1)^{+}}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k+1,j))$
$=\Delta_{n}\sum_{m=l\lor(l^{\prime}+1)}^{j-1}\int_{t_{m}}^{t_{m+1}}[a(u-t_{l})-a(u)][a_{n}((t_{l}-t_{l^{\prime}+1})^{+})-a_{n}(t_{m}-t_{l^{\prime}})]\mathrm{d}u$
@@
Hence, (75) holds with $F_{n}^{(1,3)}(s,r)$ equal to
$\frac{1}{n\Delta_{n}}\int_{s\lor r}^{n\Delta_{n}}\int_{s\lor r}^{y}[a(u-s)$ $-a(u)][a_{n}((s-r-\Delta_{n})^{+})-a_{n}(\tau_{n}(u)-s)]\mathrm{d}u\mathrm{d}y$
$+\int_{0}^{s}a(u)[a_{n}((s-r-\Delta_{n}-\tau_{n}(u))^{+})$ $-a_{n}((s-r-\Delta_{n})^{+})]\mathrm{d}u\left(1-\frac{s\lor r}{n\Delta_{n}}\right).$
In addition,
@@
$\mathbb{E}(\beta_{j,j,l}^{(1)}\chi_{j}\alpha_{k+1,j}\beta_{k,j,l^{\prime}}^{(1)})=\mathbb{E}(\beta_{k,l^{\prime},j}^{(1)}\beta_{j,l,j}^{(1)})\mathbb{E}(\alpha_{k+1,j}^{2})(1+\mathbf{1}_{k\leq j-l-1}).$
@@
$\sum_{k=l^{\prime}}^{j-1}\mathbb{E}(\varsigma_{k,j}\beta_{j,j-1,l}^{(2)})\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,j))=\Delta_{n}\sum_{m=l}^{j-l^{\prime}-1}\int_{t_{m}}^{t_{m+1}}\left[a(u-t_{l})-a(u)\right]\mathrm{d}ua_{n}(t_{m}+t_{l^{\prime}+1}).$
Therefore, (75) is fulfilled, with
@@
Finally, as above, we obtain
@@
$\Sigma_{n}^{(2,3)}(l,l^{\prime})=\frac{1}{n\Delta_{n}}\sum_{j=l^{\prime}+l^{\prime}+2}^{n}\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{j-1}\mathbb{E}(\chi_{k}\beta_{k^{\prime},j-1,l^{\prime}}^{(2)})\mathbb{E}(\alpha_{k^{\prime}+1,j}^{2})\mathbf{1}_{k^{\prime}\leq(k-l-1)\land(k+l^{\prime})}\mathbf{1}_{k\geq l+l^{\prime}+1}.$
Rearranging the sums yields
$-\Sigma_{n}^{(2,3)}(l,l^{\prime})$ $=\frac{1}{n\Delta_{n}}\sum_{k=l+l^{\prime}+1}^{n-1}\sum_{k^{\prime}=l^{\prime}}^{k-l-1}\mathrm{Leb}(\cup_{i=0}^{k^{\prime}-l^{\prime}}\mathcal{\mathcal{P}}_{A}^{n}(i,k))\sum_{j=k+1}^{n}\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k^{\prime}+1,j))$
@@
satisfying
@@
From (65),
$-\Sigma_{n}^{(2,4)}(l,l^{\prime})=$ $\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}+1}^{n-1}\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{j-1}\mathbb{E}(\chi_{k}\varsigma_{k^{\prime},j})\theta_{k,l,k^{\prime},l^{\prime}}^{j}.$
where $\theta_{k,l,k^{\prime},l^{\prime}}^{j}$ as in (71). Moreover,
$\sum_{k=l,k^{\prime}=l^{\prime}}^{j-1}\mathbb{E}(\chi_{k}\varsigma_{k^{\prime},j})\theta_{k,l,k^{\prime},l^{\prime}}^{j}$ $=\sum_{k=l,k^{\prime}=l^{\prime}}^{j-1}\mathrm{Leb}(\cup_{i=0}^{k-l}\mathcal{\mathcal{P}}_{A}^{n}(i,j))\mathrm{Leb}(\mathcal{\mathcal{P}}_{A}^{n}(k^{\prime}+1,k))\mathbf{1}_{k-1\geq k^{\prime}\geq k-(l-l^{\prime})}$
@@
Therefore, in this situation we may choose $F^{(2,4)}_{n}(s,r)$ as
$\frac{1}{n\Delta_{n}}\int_{s\lor r}^{n\Delta_{n}}\int_{s\lor r}^{y}a(u)[a_{n}((s-r)^{+})-a_{n}(\tau_{n}(u)-r-\Delta_{n})]\mathrm{d}u\mathrm{d}y$
$+[a_{n}(0)-a_{n}((s-r)^{+})]\frac{1}{n\Delta_{n}}\int_{s\lor r}^{n\Delta_{n}}\int_{s}^{y}a(u)\mathrm{d}u\mathrm{d}y,$
@@
$F^{(2,4)}(s,r)=[a(0)-a((s-r)^{+})]\int_{s}^{\infty}a(u)\mathrm{d}u+\int_{s\lor r}^{\infty}a(u)[a((s-r)^{+})-a(u-r)]\mathrm{d}u.$
@@
Finally, for all $l,l^{\prime}\geq 0$, $k\geq l$ and $k^{\prime}\geq l^{\prime}$
$\mathbb{E}[\beta_{k,j-1,l}^{(2)}\varsigma_{k^{\prime},j}]\mathbb{E}[\alpha_{k+1,j}\beta_{k^{\prime},j,l^{\prime}}^{(1)}]\mathbf{1}_{k+1\leq k^{\prime}-l^{\prime}}=\mathbb{E}[\beta_{k,j-1,l}^{(2)}\varsigma_{k^{\prime},j}]\mathbb{E}[\alpha_{k+1,j}^{2}]\mathbf{1}_{k^{\prime}+l+1\leq k\leq k^{\prime}-l^{\prime}-1}=0.$
Therefore, $\Sigma_{n}^{(3,4)}(l,l^{\prime})=0$.
We are left to show that (76) holds. Using the relation $\Sigma_{n}^{(\ell^{\prime},\ell)}(l,l^{\prime})=\Sigma_{n}^{(\ell,\ell^{\prime})}(l^{\prime},l)$ and the first part of the proof, we obtain
$\sum_{\ell,\ell^{\prime}=1}^{4}\Sigma_{n}^{(\ell,\ell^{\prime})}(\tau_{n}(s),\tau_{n}(r))\rightarrow\sum_{\ell=1}^{4}F^{(\ell,\ell)}(s,r)+\sum_{\ell=1}^{3}\sum_{\ell^{\prime}=\ell+1}^{4}\widetilde{F}^{(\ell,\ell^{\prime})}(s,r).$ (79)
Using the explicit expressions obtained above gives that
$\sum_{\ell=1}^{4}F^{(\ell,\ell)}(s,r)=\mathfrak{K}_{4}a(s\lor r)+2a(0)\int_{\rvert s-r\rvert}^{\infty}a(u)\mathrm{d}u,$
@@
$F^{(1,2)}(s,r)+F^{(1,4)}(s,r)+F^{(2,3)}(s,r)=\sigma_{3}(s,r),$
and
$F^{(1,3)}(s,r)+F^{(2,4)}(s,r)=\sigma_{2}(s,r)-a(0)\int_{|s-r|}^{\infty}a(u)\,\mathrm{d}u.$
Plugging in the previous equations into (79) completes the proof.∎